OpenStudy (anonymous):

Trig problem.. find all solutions in the interval [0,360°) 10 cos^2x+11cosx+3=0 plz show steps thanx!

OpenStudy (anonymous):

can you factor the equation?

OpenStudy (anonymous):

umm yes i think so

OpenStudy (anonymous):

ok then do it

OpenStudy (anonymous):

I think I factored the left side.... is this correct? 5cosx+3=0 5cosx-3=0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

honestly?? well how else do I factor out 10cos besides using 5? am I labeling wrong or something?

OpenStudy (anonymous):

let u = cosx can you factor 10u^2 + 11u + 3?

OpenStudy (anonymous):

use AC method

OpenStudy (anonymous):

you're looking for two integers that when multiplied gives 30 and when added gives 11

OpenStudy (anonymous):

found those two integers?

OpenStudy (anonymous):

No i'm totally lost on this i think i'll just do a similar problem i'm really dumb at math thnx for working w/ me though.

OpenStudy (anonymous):

do 5 and 6 work?

OpenStudy (anonymous):

Ohhhh no see what you were doing but I'm doing this. I meant to input that x=____

OpenStudy (anonymous):

idk why i didn't think to snip it earlier sorry for the confusion

OpenStudy (anonymous):

well, you find x by factoring first, which requires you to do the ac method

OpenStudy (anonymous):

can you now factor 10u^2 + 5u + 6u + 3?

OpenStudy (anonymous):

"u" is theta?

OpenStudy (anonymous):

no, u = cosx. The reason I let u = cosx is because it's easy to work with that way

OpenStudy (anonymous):

you could have done 10cos^2(x) + 5cos(x) + 6cos(x) + 3 but it's alittle bit harder to see

OpenStudy (anonymous):

so instead, let u = cosx, so you have 10u^2 + 5u + 6u + 3

OpenStudy (anonymous):

hmmm?? I never tried it this way before but is it (5u+3) (2u+1)

OpenStudy (anonymous):

yes. so can you solve (5u+3) (2u+1) = 0 ?

OpenStudy (anonymous):

ummm -3/5, -1/2 ???

OpenStudy (anonymous):

yes, u = -3/5 or u = -1/2 but u = cos(x), so cos(x) = -3/5 or cos(x) = -1/2 can you solve for x?

OpenStudy (anonymous):

cos(x)=1/2 x=2pi/3=120° & 4pi/3=240°

OpenStudy (anonymous):

it's cos(x) = -1/2

OpenStudy (anonymous):

I used the neg sign just forgot to type it but... You know? I think I have it now!!!!! thanxx :) for the help!

OpenStudy (anonymous):

no prob